# Overview of Speaker Modeling and Its Applications: From the Lens of Deep Speaker Representation Learning

### v1

@@ Abstract
Speaker individuality information is among the most critical elements within speech signals. By thoroughly and accurately modeling this information, it can be utilized in various intelligent speech applications, such as speaker recognition, speaker diarization, speech synthesis, and target speaker extraction. In this article, we aim to present, from a unique perspective, the developmental history, paradigm shifts, and application domains of speaker modeling technologies within the context of deep representation learning framework. This review is designed to provide a clear reference for researchers in the speaker modeling field, as well as for those who wish to apply speaker modeling techniques to specific downstream tasks.
@@
Speaker embedding learning, overview, speaker modeling
@@ I-D Organization of This Overview
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hiddendraw, rounded corners, align=left, minimum width=2.5em, inner xsep=4pt, inner ysep=0pt, , where level=1text width=14.5em,align=center,font=, where level=2text width=19.5em,align=left,font=, [Speaker Modeling [Fundamentals [Sec. I: Introduction] [Sec. II: Terminology and Task Formulation] [Sec. III: Speaker Encoders] ] [Sec. IV: Paradigm Shift [Sec. IV-A: Supervised to Self-supervised] [Sec. IV-B: Semi-superivised Training] [Sec. IV-C: Leverage Pretrained Models] [Sec. IV-D: Unimodal to Multi-modal] ] [Sec. V-A: Robustness & Efficiency [Sec. V-A1: Text Dependency] [Sec. V-A2: Cross-language Recognition] [Sec. V-A3: Channel Variability] [Sec. V-A4: Intra-speaker Variability] [Sec. V-A5: Duration Variability] [Sec. V-A6: Learning from Noisy Labels] [Sec. V-B1: Model Quantization] [Sec. V-B2: Knowledege Distillation] [Sec. V-B3: Efficient Architectures] [Sec. V-B4: Efficient Policy] ] [Sec. VI: Explainablity [Sec. VI-A: Probing Tasks] [Sec. VI-B: Component Ablation] [Sec. VI-C: Relevance analysis] [Sec. VI-D: Visualization] [Sec. VI-E: Explainable Network Design] ] [Sec. 3: Applications & Strategies [Sec. VII-A: Integration Strategies] [Sec. VII-B: Speaker Diarization] [Sec. VII-C: Speech Recognition] [Sec. VII-D: Speech Generation] [Sec. VII-E: Target speaker front-ends] ] [Sec. VIII: Dataset and Tools [Sec. VIII-A: Dataset] [Sec. VIII-B: Open-source Toolkits] ] [Sec. IX: Trends [Sec. IX-A: Privacy and Ethics] [Sec. IX-B: Attacks and Defenses] [Sec. IX-C: Controlable Speaker Generation] [Sec. IX-D: Disentanglement Learning] ] [Sec. X: Conclusion] ]
@@ II Formulation of The Deep Speaker Representation Learning Problem
Speaker representation learning is the process of extracting compact and discriminative representations from given speech signals. These representations aim to capture the unique acoustic characteristics of speakers and maintain consistency across different languages, content, and environmental conditions. Mathematically, this can be defined as a mapping function $f$ that maps a speech signal $O={o_{1},o_{2},\ldots,o_{T}}$ of length $T$ to a fixed-length vector $\mathbf{v}$, which represents the speaker’s identity:
$\mathbf{v}=f(O;\Theta)$ (1)
@@ III Speaker Encoders
III-A Frame-Level Encoders
Represented by d-vector [49], the process of speaker representation learning optimized at the frame level is illustrated in Figure 3. The neural network takes frame-level features as input and optimizes the cross-entropy (CE) loss function at this granularity. Once the network training is complete, for an input speech feature comprising $T$ frames, denoted by $\mathbf{O}={\mathbf{o}_{1},\cdots,\mathbf{o}_{T}}$, we extract frame-level outputs from the hidden layers close to the output layer, represented by $\mathbf{F}={\mathbf{f}_{1},\cdots,\mathbf{f}_{T}}$. Then, an aggregation operation is applied to transform the sequence of frame-level representations into a sentence-level speaker representation. The most common method of aggregation is averaging:
@@ III-A Frame-Level Encoders
III-B Segment-Level Encoders
@@ IV-A2 Self-Supervised Methods
[…]
where $d(\cdot,\cdot)$ can be any distance metric formula, and the loss function does not have to be as simple as the one above. However, the ultimate goal is to make the distance between positive pairs smaller and the distance between negative pairs larger. For instance, Jati et al. [156] first measured the distance between two segments using $L_{1}$ distance and then used a binary classification loss to distinguish the positive and negative pair. Ravanelli et al. [157] maximize the mutual information (MI) between positive pairs and minimize it between negative pairs. To make the model more robust to the channel variation, the authors [158] added different data augmentations to the sampled segments from the same utterance and added an extra loss function to constrain the distance for the positive pairs. Besides, the authors in [159] leveraged extra adversarial training loss to help the system more robust to the noise. To enrich the diversity of the negative pair data in one batch, Xia et al. [160] employed a buffer to store the speaker embedding samples from the previous batches.
@@ VII-A Integration Strategies
Generally speaking, strategies for integrating speaker information into other tasks can be mainly divided into two categories.
@@ VII-E Target Speaker Frond-End processing
Speech front-end processing algorithms have a wide range of applications, including tasks such as speech enhancement, voice activity detection, and speech separation. Sometimes, we want to focus specifically on front-end processing algorithms for a particular speaker, which has led to the development of a series of speaker-dependent front-end processing algorithms such as personal voice activity detection [334, 335], target speaker enhancement, and target speaker extraction [19]. These methods often specify the target speaker by introducing additional speaker encoding, which can be either pretrained [336, 337, 338, 339] or jointly trained [340, 341, 342].
@@ VIII Benchmark Datasets and Open-Source Toolkits
In recent years, we have seen the release of several new datasets such as VoxCeleb [162], CNCeleb [254], and 3D-Speaker [343]. Alongside these, numerous open-source tools [344, 345, 346, 347, 348] dedicated to speaker representation learning have emerged, including the WeSpeaker project initiated by ourselves [348]. In this section, we will provide a summarization of these popular datasets and open-source toolkits.
@@ VIII-B Open-Source Toolkits
Easy-to-use open-source tools are key to advancing the development of related fields. We list the open-source toolkits related to speaker recognition in Table II. Firstly, Kaldi, a speech tool popular for over a decade, offers many recipes for the speaker recognition task. Nowadays, many general-purpose speech open-source tools have also incorporated recipes related to speaker recognition, such as Espnet [381], Speechbrain [346], and NeMo. Additionally, there are some specialized open-source tools for speaker recognition, like the early ASV-Subtools [347] which were compatible with Kaldi’s data format. Later speaker recognition tools have largely abandoned the Kaldi format, making the entire framework easier for researchers to modify and add their own modules. Moreover, self-supervised speaker recognition has received widespread attention in recent years, and some open-source frameworks, such as 3D-speaker [382] and Wespeaker [348], have also supported this. Furthermore, Wespeaker and NeMo also offer rapid deployment features to assist the industry in quickly deploying models.
@@
Toolkit Speaker-task specific SSL Pretrained Speech Models Deployment Kaldi [55] No No No No VoxCeleb_Trainer [162] Yes No No No ASV-Subtools [347] Yes No No No SpeechBrain [346] No No No No NeMo [^4] https://docs.nvidia.com/nemo-framework/user-guide/latest/nemotoolkit/asr/speaker_recognition/intro.html No No No Yes Espnet [381] No No Yes No 3D-Speaker [382] Yes Yes No No Wespeaker [348] Yes Yes No Yes
@@ IX-A Privacy Protection and Ethical Issues
IX-B Attacks and Defenses
@@
IX-C Controlable Speaker Generation

### v2

@@ Abstract
Speaker individuality information is among the most critical elements within speech signals. By thoroughly and accurately modeling this information, it can be utilized in various intelligent speech applications, such as speaker recognition, speaker diarization, speech synthesis, and target speaker extraction. In this overview, we present a comprehensive review of neural approaches to speaker representation learning from both theoretical and practical perspectives. Theoretically, we discuss speaker encoders ranging from supervised to self-supervised learning algorithms, standalone models to large pretrained models, pure speaker embedding learning to joint optimization with downstream tasks, and efforts toward interpretability. Practically, we systematically examine approaches for robustness and effectiveness, introduce and compare various open-source toolkits in the field. Through the systematic and comprehensive review of the relevant literature, research activities, and resources, we provide a clear reference for researchers in the speaker characterization and modeling field, as well as for those who wish to apply speaker modeling techniques to specific downstream tasks.
@@
Speaker embedding learning, speaker modeling, speaker recognition, overview, survey
@@ I-D Organization of This Overview
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hiddendraw, rounded corners, align=left, minimum width=2.5em, inner xsep=4pt, inner ysep=0pt, , where level=1text width=14.5em,align=center,font=, where level=2text width=19.5em,align=left,font=, [Speaker Modeling [Fundamentals [Sec. I: Introduction] [Sec. II: Terminology and Task Formulation] [Sec. III: Speaker Encoders] ] [Sec. IV: Paradigm Shift [Sec. IV-A: Supervised to Self-supervised] [Sec. IV-B: Semi-superivised Training] [Sec. IV-C: Leverage Pretrained Models] [Sec. IV-D: Unimodal to Multi-modal] ] [Sec. V-A: Robustness & Efficiency [Sec. V-A1: Text Dependency] [Sec. V-A2: Cross-language Recognition] [Sec. V-A3: Channel Variability] [Sec. V-A4: Intra-speaker Variability] [Sec. V-A5: Duration Variability] [Sec. V-A6: Learning from Noisy Labels] [Sec. V-B1: Model Quantization] [Sec. V-B2: Knowledege Distillation] [Sec. V-B3: Efficient Architectures] [Sec. V-B4: Efficient Policy] ] [Sec. VI: Explainablity [Sec. VI-A: Probing Tasks] [Sec. VI-B: Component Ablation] [Sec. VI-C: Relevance analysis] [Sec. VI-D: Visualization] [Sec. VI-E: Explainable Network Design] ] [Sec. 3: Applications & Strategies [Sec. VII-A: Integration Strategies] [Sec. VII-B: Speaker Diarization] [Sec. VII-C: Speech Recognition] [Sec. VII-D: Speech Generation] [Sec. VII-E: Target Speaker Front-Ends] ] [Sec. VIII: Dataset and Tools [Sec. VIII-A: Dataset] [Sec. VIII-B: Open-Source Toolkits] ] [Sec. IX: Trends [Sec. IX-A: Privacy and Ethics] [Sec. IX-B: Large-Scale SSL] [Sec. IX-C: Speaker-Text Cross Modality] [Sec. IX-D: Attacks and Defenses] [Sec. IX-E: Controlable Speaker Generation] [Sec. IX-F: Disentanglement Learning] ] [Sec. X: Conclusion] ]
@@ II Formulation of The Deep Speaker Representation Learning Problem
Speaker representation learning is the process of extracting compact and discriminative representations from given speech signals. These representations aim to capture the unique acoustic characteristics of speakers and maintain consistency across different languages, content, and environmental conditions. Mathematically, this can be defined as a mapping function $f$ that maps a speech signal $X=\{x_{1},x_{2},\ldots,x_{T}\}\in\mathbb{R}^{T}$ of length $T$ to a fixed-length vector $\mathbf{v}$, which represents the speaker’s identity:
$\mathbf{v}=f(X;\Theta)$ (1)
@@ III Speaker Encoders
III-A Speaker Encoders Optimized at Frame-Level
Represented by d-vector [49], the process of speaker representation learning optimized at the frame level is illustrated in Figure 3. The neural network takes frame-level features as input and optimizes the cross-entropy (CE) loss function at this granularity. Once the network training is complete, for an input speech feature comprising $T$ frames, denoted by $\mathbf{O}=\{\mathbf{o}_{1},\cdots,\mathbf{o}_{T}\}\in\mathbb{R}^{T\times D}$, we extract frame-level outputs from the hidden layers close to the output layer, represented by $\mathbf{F}=\{\mathbf{f}_{1},\cdots,\mathbf{f}_{T}\}\in\mathbb{R}^{T\times D^{\prime}}$. Then, an aggregation operation is applied to transform the sequence of frame-level representations into a sentence-level speaker representation. The most common method of aggregation is averaging:
@@ III-A Speaker Encoders Optimized at Frame-Level
III-B Speaker Encoders Optimized at Segment-Level
@@ IV-A2 Self-Supervised Methods
$\mathcal{L_{\textit{con}}}=\sum_{N^{+}}\operatorname{d}\left(x_{i},x_{i}^{+}\right)-\sum_{N^{-}}\operatorname{d}\left(x_{i},x_{i}^{-}\right)$ (4)
where $\operatorname{d}(\cdot,\cdot)$ can be any distance metric formula, and the loss function does not have to be as simple as the one above. However, the ultimate goal is to make the distance between positive pairs smaller and the distance between negative pairs larger. For instance, Jati et al. [156] first measured the distance between two segments using $L_{1}$ distance and then used a binary classification loss to distinguish the positive and negative pair. Ravanelli et al. [157] maximize the mutual information (MI) between positive pairs and minimize it between negative pairs. To make the model more robust to the channel variation, the authors [158] added different data augmentations to the sampled segments from the same utterance and added an extra loss function to constrain the distance for the positive pairs. Besides, the authors in [159] leveraged extra adversarial training loss to help the system more robust to the noise. To enrich the diversity of the negative pair data in one batch, Xia et al. [160] employed a buffer to store the speaker embedding samples from the previous batches.
@@ VII-A Integration Strategies
Generally speaking, strategies for integrating speaker information into other tasks can be mainly divided into two categories: off-the-the usage and joint optimization. Regardless of the strategy, we need specific fusion methods to integrate speaker information into other tasks. Additionally, both off-the-shelf usage and joint optimization have their advantages and disadvantages, and the choice between them should be based on specific needs during implementation.
@@ VII-E Target Speaker Frond-End processing
Speech front-end processing algorithms have a wide range of applications, including tasks such as speech enhancement, voice activity detection, and speech separation. Sometimes, we want to focus specifically on front-end processing algorithms for a particular speaker, which has led to the development of a series of speaker-dependent front-end processing algorithms such as personal voice activity detection [334, 335], target speaker enhancement, and target speaker extraction [19, 336]. These methods often specify the target speaker by introducing additional speaker encoding, which can be either pretrained [337, 338, 339, 340] or jointly trained [341, 342, 343].
@@ VIII Benchmark Datasets and Open-Source Toolkits
In recent years, we have seen the release of several new datasets such as VoxCeleb [162], CNCeleb [254], and 3D-Speaker [344]. Alongside these, numerous open-source tools [345, 346, 347, 348, 349] dedicated to speaker representation learning have emerged, including the WeSpeaker project initiated by ourselves [349, 350]. In this section, we will provide a summarization of these popular datasets and open-source toolkits.
@@ VIII-B Open-Source Toolkits
Accessible and user-friendly open-source tools are pivotal for advancing research in the field of speaker recognition. Table II enumerates the relevant open-source toolkits. Kaldi, a speech processing toolkit that has remained popular for over a decade, provides numerous recipes for speaker recognition tasks. Additionally, many general-purpose speech platforms have integrated speaker recognition functionalities, including Espnet [383], Speechbrain [347], and NeMo. Specialized open-source tools for speaker recognition, such as ASV-Subtools [348], VoxCeleb_Trainer [162], 3D-speaker [384], and Wespeaker [349], are also available. These toolkits are implemented in PyTorch and are easily modifiable, facilitating the extension of their frameworks by researchers. Furthermore, self-supervised speaker recognition has garnered significant attention in recent years. Open-source frameworks like 3D-speaker [384] and Wespeaker [350] support such functionalities. To leverage the capabilities of large pretrained speech models, Espnet-Spk [383] and Wespeaker [350] enable the integration of pretrained models such as WavLM and Whisper. To bridge the gap between research and real-world applications, Wespeaker and NeMo offer rapid deployment features, enabling swift implementation of models within industrial settings.
@@
Toolkit Speaker-task specific SSL Pretrained Speech Models Deployment Kaldi [55] No No No No VoxCeleb_Trainer [162] Yes No No No ASV-Subtools [348] Yes No No Yes SpeechBrain [347] No No No No NeMo [^4] https://docs.nvidia.com/nemo-framework/user-guide/latest/nemotoolkit/asr/speaker_recognition/intro.html No No No Yes Espnet [383] No No Yes No 3D-Speaker [384] Yes Yes No No Wespeaker [349] Yes Yes Yes Yes
@@ IX-A Privacy Protection and Ethical Issues
IX-B Large Scale Self-supervised Representation Learning
Research from many other fields has already demonstrated that large-scale pretraining can enhance model performance, such as Whisper [391] in speech recognition and BaseTTS [392] in the TTS field. As we mentioned in the previous section, data with speaker labels involves privacy issues, which may prevent us from using large amounts of labeled data to learn speaker representations. Furthermore, the pretrained model introduced in section IV-C was not specifically designed for speaker representation learning during training. Additionally, the speaker self-supervised representation learning described in section IV-A2 has only been validated on datasets at the scale of Voxceleb. Moreover, the research [135] points out that the related algorithms do not bring performance improvements when scaled to larger datasets. Therefore, further research is needed if we want to use larger-scale unlabeled data for speaker representation learning.
IX-C The Cross-modality Learning between Text and Speaker Representation
Recently, large language models (LLMs) [393, 394, 395] have garnered widespread attention from researchers. Beyond modeling text itself, many researchers have found that applying LLMs to various cross-modal fields can also be highly beneficial. In cross-modal tasks, LLMs [396, 397] are often used as an intermediary to connect different modalities. Similarly, it’s crucial to explore how large language models can understand speaker representations, that is, converting speaker representations into textual descriptions. Conversely, it is also important to study how to convert textual descriptions of speaker characteristics into speaker representations. Regarding the understanding of speaker characteristics using text, ByteDance has proposed the SALMONN [398] model, which can use text to output speaker characteristics in speech. Similarly, there are other studies that generate speaker characteristics [399, 400, 401, 402] based on text descriptions, but these studies are still in their early stages, and the performance is not very good. Further research is needed in the future.
IX-D Attacks and Defenses
@@
IX-E Controlable Speaker Generation